OpenStudy (anonymous):

Solve: I x+4 I=7 Answer A. x = 3, -11 B. x = -3, 11 C. x = -3, -11 D. x = 3, 11

OpenStudy (anonymous):

Choice 4

Directrix (directrix):

I x+4 I=7 means (x + 4) = 7 OR (x + 4) = -7 x = 3 OR x = -7 -4 = -11 The two solutions are 3 and -11. ---> D. x = 3, 11